# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v3

@@ 4.1 Experimental Setup
LLM Models. We compare LLMs that use the Transformer model architecture (Vaswani et al., 2017). We compare models that use the same tokenizer because the highest activation correlation pairing relies on comparing two activations using the same tokens. We compare the residual streams of Pythia-70m, which has 6 layers and 512 dimensions, to Pythia-160m, which has 12 layers and 768 dimensions (Biderman et al., 2023). We compare the residual streams of Gemma-1-2B, which has 18 layers and 2048 dimensions, to Gemma-2-2B, which has 26 layers and 2304 dimensions (Team et al., 2024a; b).  [^2] Gemma-2 introduces several architectural improvements over Gemma-1 (Team et al., 2024a; b). We compare SAEs trained on the same Tinystories model (Eldan & Li, 2023) in Appendix §E, and show results for across Llama 3 and 3.1 models (Dubey et al., 2024; AI, 2024) in Appendix §C.3. In total, we compare 5 model pairs and find notable similarities across their SAEs. The list of model pairs is given in Appendix A.
SAE Models. For Pythia, we use pretrained SAEs with 32768 feature neurons trained on residual stream layers from Eleuther (EleutherAI, 2023). For Gemma, we use pretrained SAEs with 16384 feature neurons trained on residual stream layers (Lieberum et al., 2024) We access pretrained Gemma-1-2B SAEs for layers 0, 6, 10, 12, and 17, and Gemma-2-2B SAEs for layers 0 to 26, through the SAELens library (Bloom & Chanin, 2024).
In Appendix §C, we compare SAEs for Gemma-2-2B vs Gemma-2-9B, base vs fined tuned versions for Gemma-2-9B vs Gemma-2-9B-Instruct and Gemma-1-2B vs Gemma-1-2B-Instruct (at Layer 12), and for Llama 3-8B and Llama 3.1-8B (at Layer 25). The list of all models used in experiments is provided in Section §A. In Appendix §D, we assess a baseline comparison that compares Pythia-70m SAEs with SAEs trained on a randomized Pythia-70m. In Appendix §E, we compare SAEs trained within the same model. Lastly, in Appendix §F, we compare SAEs across Pythia-70m and Pythia-160m as we vary SAE widths.
Datasets. We obtain SAE activations using OpenWebText, a dataset that scraps internet content (Gokaslan & Cohen, 2019). We use 100 samples with a max sequence length of 300 for Pythia for a total of 30k tokens, and we use 150 samples with a max sequence length of 150 for Gemma for a total of 22.5k tokens. Top activating dataset tokens were also obtained using OpenWebText samples. In Appendix §G, we vary the dataset and number of tokens used to obtain activations to show that our results do not drastically change when these factors are varied.
Computing Resources. We run experiments on an A100 GPU.
@@ 4.2 SAE Space Similarity
In summary, for almost all layer pairs after layer 0, we find the p-value of SAE experiments for SVCCA, which measures rotationally-invariant alignment, to be low; all of them are between 0.00 to 0.01 using 100 samples. [^3] Only 100 samples were used as we found the variance to be low for the null distribution, and that the mean was similar for using 100 vs 1k samples. While RSA also passes many p-value tests, their values are much lower, suggesting a relatively weaker similarity in terms of pairwise differences. It is possible that many features are not represented in just one layer; hence, a layer from one model can have similarities to multiple layers in another model. We also note that the first layer of every model, layer 0, has a very high p-value when compared to every other layer, including when comparing layer 0 from another model. This may be because layer 0 contains few discernible, meaningful, and comparable features.
As shown in Figure 24 for Pythia and Figure 25 for Gemma in Appendix M, we also find that mean of the highest activation correlations does not always correlate with the global similarity metric scores. For instance, a subset of feature pairings with a moderately high mean activation correlation (eg. 0.6) may yield a low SVCCA score (eg. 0.03).
Because LLMs, even with the same architecture, can learn many different features, we do not expect entire feature spaces to be highly similar; instead, we hypothesize there may be highly similar subspaces. In other words, there are many similar features universally found across SAEs, but not every feature learned by different SAEs is the same. Our experimental results support this hypothesis.
Pythia-70m vs Pythia-160m . In Figure 20, we compare Pythia-70m, which has 6 layers, to Pythia-160m, which has 12 layers. When compared to Figures 26 and 27 in Appendix M, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. In particular, Figure 20 shows that Layers 2 and 3, which are middle layer of Pythia-70m, are more similar by SVCCA and RSA scores to middle layers 4 to 7 of Pythia-160m compared to other layers. Figure 20 also shows that Layer 5, the last residual stream layer of Pythia-70m, is more similar to later layers of Pythia-160m than to other layers. The number of features after filtering non-concept features and many-to-1 features are given in Tables 5 and 6 in Appendix M.
@@
Rather than just finding that middle layers are more similar to middle layers, we also find that they are the best at representing concepts. This is consistent with work by (Rimsky et al., 2024), which find the best steering vectors in middle layers for steering behaviors such as sycophancy. Other papers find that middle layers tend to provide better directions for eliminating refusal behavior (Arditi et al., 2024) and work well for representing goals and finding steering vectors in policy networks (Mini et al., 2023). In fact, we find that early layers do not represent universal features well, perhaps due to lacking content, while later layers do not represent them as well as middle layers, perhaps due to being too specific/complex.
@@ 4.3 Similarity of Semantically Matched SAE Feature Spaces
We find that for all layers and for all concept categories, Test 2 described in §3 is passed. Thus, we only report specific results for Test 1 in Tables 1 and 2. Overall, in both Pythia and Gemma models and for many concept categories, we find that semantic subspaces are more similar to one another than non-semantic subspaces.
Pythia-70m vs Pythia-160m. We compare every layer of Pythia-70m to every layer of Pythia-160m for several concept categories. While many layer pairs have similar semantic subspaces. Middle layers appear to have the semantic subspaces with the highest similarity. Table 1 demonstrates one example of this by comparing the SVCCA score for layer 3 to layer 5 of Pythia-160m, which shows high similarity for semantic subspaces across models, as the concept categories all pass Test 1, having p-values below 0.05. We show the scores for other layer pairs and categories in Figures 30 and 32 in Appendix M, which include RSA scores in Table 9.
@@ 5 Related Work
Superposition and Sparse Autoencoders. Superposition is a phenomenon in which a model, in response to the issue of having to represent more features than it has parameters, learns feature representations distributed across many parameters (Elhage et al., 2022). This causes its parameters, or neurons, to be polysemantic, which means that each neuron is involved in representing more than one feature. This leads to issues for interpretability, which aims to disentangle and cleanly identify features in models in order to manipulate them, such as by steering a model away from deception (Templeton et al., 2024). To address the issue of polysemanticity, Sparse Autoencoders (SAEs) have been applied to disentangle an LLM’s polysemantic neuron activations into monosemantic "feature neurons", which are encouraged to represent an isolated concept (Makhzani & Frey, 2013; Cunningham et al., 2023; Gao et al., 2024; Rajamanoharan et al., 2024a; b). Features interactions in terms of circuits have also been studied (Marks et al., 2024).
Feature Universality. To the best of our knowledge, only Bricken et al. (2023) has done a quantitative study on individual SAE feature similarity for two 1-layer toy models, finding that individual SAE feature correlations are stronger than individual LLM neuron correlations; however, this study did not analyze the global properties of feature spaces. Gurnee et al. (2024) found evidence of universal neurons across language models by applying pairwise correlation metrics, and taxonimized families of neurons based on patterns of downstream, functional effects. O’Neill et al. (2024) discover "feature families", which represent related hierarchical concepts, in SAEs across two different datasets.
Of note is that our paper seeks to measure the extent of universality in terms of analogous features and representational universality, as defined in previous work (Olah et al., 2020; Gurnee et al., 2024), but we do not measure true features, which are exact, atomic hypothesized ground truth features that models may converge towards capturing. This is explained further in Appendix §H. Recent work has shown that as vision and language models are trained with more parameters and with better methods, their representational spaces converge towards more similar representations (Huh et al., 2024). However, we do not claim nor attempt to study in this paper if models converge towards finding the same features.
@@
Representational Space Similarity. Previous work has studied neuron activation spaces by utilizing metrics to compare the geometric similarities of representational spaces (Raghu et al., 2017; Wang et al., 2018b; Kornblith et al., 2019; Klabunde et al., 2024; 2023; Kriegeskorte et al., 2008; Sucholutsky et al., 2023). It was found that even models with different architectures may share similar representation spaces, hinting at feature universality. However, these techniques have yet to be applied to the feature spaces of sparse autoencoders trained on LLMs. Moreover, these techniques compare the similarity of paired input activations. Our work differs as it compares the similarity of paired feature weights. Essentially, previous works do not show that one can match LLMs by features that both LLMs find in common.
Feature Manifolds. In real world data, features may lie on a manifold, where nearby features respond to similar data. (Olah & Batson, 2023; Bricken et al., 2023). Previous work has discovered the existence of features corresponding to months and days of the week arranged in a circular structure in SAE feature spaces across models (Engels et al., 2024a). Studying how feature arrangements generalize across models may shed light on how features lie on a manifold.
@@
Mechanistic Interpretability. Recent work has made notable progress in neuron interpretation (Foote et al., 2023; Garde et al., 2023) and interpreting the interactions between Attention Heads and MLPs (Neo et al., 2024). Other work in mechanistic interpretability has traditionally focused on ’circuits’-style analysis (Elhage et al., 2021), as well as the use of steering vectors to control model behaviors at inference using identified representations (Zou et al., 2023; Turner et al., 2023; Bereska & Gavves, 2024). These approaches have been combined with SAEs to steer models in the more interpretable SAE feature space (Nanda & Conmy, 2024). Steering vectors have been found across models; for instance, Arditi et al. (2024) found "Refusal Vectors" that steered a model to refuse harmful instructions across 13 different models, including Gemma and Llama-2.
@@
AI Safety. Advanced AI systems may not be well aligned with the values of their designers (Ngo et al., 2022), due to phenomena such as goal misgeneralization (Shah et al., 2022; Langosco et al., 2022), deceptive alignment (Hubinger et al., 2024; 2021; Greenblatt et al., 2024), sycophancy (Sharma et al., 2023), and reward tampering (Denison et al., 2024). SAEs may be able to find AI safety-related features across models; using SAEs to reliably find universal capabilities, especially in combination with control methods such as steering Chalnev et al. (2024), may help provide solutions to the above pheonomena.
@@ 6 Conclusion
In this study, we address the underexplored domain of feature universality in sparse autoencoders (SAEs) trained on LLMs with multiple layers and across multiple model pairs. To achieve these insights, we developed a novel methodology to pair features with high mean activation correlations, and then assess their global similarity of the resulting subspaces using SVCCA and RSA. Our findings reveal a high degree of similarity in SAE feature subspaces across various models. Furthermore, our research reveals that subspaces of features associated with specific semantic concepts, such as calendar or people tokens, demonstrate remarkably high similarity across different models. This suggests that certain semantic feature subspaces are universally encoded across varied LLM architectures. However, our results also show that there are parts of SAE feature spaces which do not match, and may suggest that either certain features are not captured universally, or that SAEs have limitations in capturing universal features. We believe that future work can benefit from these discoveries to find better ways on improving methods to capture feature universality. Our work enhances understanding of how LLMs represent information at a fundamental level, opening up new avenues for further research into model interpretability and the transferability of learned features.
@@
We perform analysis only on a sample of SAEs for similar LLMs that use the same tokenizer, and focus on SAEs with the same or similar number of features. As we pair featuers by activation correlation, we do not perform analysis on models that use different tokenizers. Although, for some examples, we obtain good results when pairing featuers only by semantic similarity, these results were not consistent when applied to other examples, perhaps due to the lower pairing quality, as the representational similarity metrics depend on the pairing choice. Future work can analyze models whose representations do not depend as much on the tokenizer choice (team et al., 2024).
Our study is also limited by the inherent challenges in interpreting and comparing high-dimensional feature spaces. The methods we used to measure similarity, while robust, may not capture all nuances of feature interactions and representations within the SAEs. Additionally, our analysis is constrained by the computational resources available, which may have limited the scale and depth of our comparisons. The generalizability of our findings to SAEs trained on more diverse datasets or specialized domain-specific LLMs remains an open question. Furthermore, the static nature of our analysis does not account for potential temporal dynamics in feature representation that might occur during the training process of both LLMs and SAEs. Lastly, functional similarity experiments were not examined.
Impact Statement
Our study investigates feature universality in large language models (LLMs) using sparse autoencoders (SAEs) to identify common feature spaces across models. This research contributes to safety research by providing insights into how different LLMs encode safety-related concepts, which has several ethical and societal implications. By identifying universal feature spaces, our approach may aid in the detection of harmful or deceptive model behaviors across different architectures. Understanding how models encode knowledge could lead to improved transparency and control mechanisms, reducing risks associated with goal misgeneralization and deceptive alignment in AI systems. Additionally, identifying shared feature structures can inform bias mitigation strategies, as universal features could either amplify or help detect biases present across multiple models.
@@
Michael conceived the study and designed the experiments with help from Fazl. Michael performed the experiments with help from Austin and led the writing of the manuscript. Philp, Ashkan and David provided advice throughout the project. Fazl advised this work and helped significantly with the write-up of the paper.

### v4

@@ 4.1 Experimental Setup
We run experiments on an A100 GPU.
LLM Models. We compare LLMs that use the Transformer model architecture (Vaswani et al., 2017). We compare models that use the same tokenizer because the highest activation correlation pairing relies on comparing two activations using the same tokens. In the main text, we compare the following residual stream layers pairs: (1) Pythia-70m, which has 6 layers and 512 dimensions, to Pythia-160m, which has 12 layers and 768 hidden dimensions (Biderman et al., 2023), (2) Gemma-1-2B, which has 18 layers and 2048 dimensions, to Gemma-2-2B, which has 26 layers and 2304 dimensions (Team et al., 2024a, b), (3) Gemma-2-2B to Gemma-2-9B, which has 42 layers and 3584 dimensions, and (4) Llama 3-8B-Instruct and Llama 3.1-8B, which both have 32 layers and 4096 dimensions. Appendix §A lists all 8 model pairs we compare.
SAE Models. For Pythia models, we use SAEs with 32768 feature neurons trained on residual stream layers from Eleuther (EleutherAI, 2023). For the Gemma models, we use SAEs with 16384 feature neurons trained on residual stream layers (Lieberum et al., 2024). The Llama 3 SAE has a width of 65536 features and uses Gated ReLU (Rajamanoharan et al., 2024a), while the Llama 3.1 SAE has a width of 32768 features and uses JumpReLU (Rajamanoharan et al., 2024b). We access pretrained Gemma and Llama SAEs through the SAELens library (Bloom and Chanin, 2024), which only hosted one SAE for Llama 3 at the time of our work.
In Appendix §C, we compare SAEs for base vs fined tuned versions for Gemma-2-9B vs Gemma-2-9B-Instruct and Gemma-1-2B vs Gemma-1-2B-Instruct (at Layer 12). In Appendix §D, we assess a baseline comparison that compares Pythia-70m SAEs with SAEs trained on a randomized Pythia-70m. In Appendix §E, we compare SAEs trained within the same model. In Appendix §F, we compare SAEs across Pythia-70m and Pythia-160m as we vary SAE widths.
Datasets. We obtain SAE activations using OpenWebText, a dataset that scraps internet content (Gokaslan and Cohen, 2019). We use 100 samples with a max sequence length of 300 for Pythia for a total of 30k tokens, and we use 150 samples with a max sequence length of 150 for Gemma for a total of 22.5k tokens. Top activating dataset tokens were also obtained using OpenWebText samples. In Appendix §G, we vary the dataset and number of tokens used to obtain activations to show that our results do not drastically change when these factors are varied.
@@ 4.2 SAE Space Similarity
In summary, for almost all layer pairs after layer 0, we find the p-value of SAE experiments for SVCCA, which measures rotationally-invariant alignment, to be low; most are between 0.00 to 0.01 using 100 samples. [^3] Only 100 samples were used as we found the variance to be low for the null distribution, and that the mean was similar for using 100 vs 1k samples. While RSA also passes many p-value tests, their values are much lower, suggesting a relatively weaker similarity in terms of pairwise differences. It is possible that many features are not represented in just one layer; hence, a layer from one model can have similarities to multiple layers in another model. We also note that the first layer of every model, layer 0, has a very high p-value when compared to every other layer, including when comparing layer 0 from another model. This may be because layer 0 contains few discernible, meaningful, and comparable features. We find that middle layer pairs have the highest scores. On the contrary, early layers have much lower scores, perhaps due to lacking content, while later layers also have scores lower than middle layers, perhaps due to being too specific/complex.
As shown in Figure 24 for Pythia and Figure 25 for Gemma-1 vs Gemma-2 in Appendix M, we also find that mean of the highest (local) activation correlations does not always correlate with the global similarity measure scores. For instance, a subset of feature pairings with a moderately high mean activation correlation (eg. 0.6) may yield a low SVCCA score (eg. 0.03).
Because LLMs, even with the same architecture, can learn many different features, we do not expect entire feature spaces to be highly similar; instead, we hypothesize there may be highly similar subspaces. In other words, there are many similar features universally found across SAEs, but not every feature learned by different SAEs is the same. Our results support this hypothesis. For all experiments in this section, approximately 10-30% (and typically 20%) of feature pairs are kept after filtering both non-concept and many-1 pairings.
Pythia-70m vs Pythia-160m. In Figure 2, we compare Pythia-70m, which has 6 layers, to Pythia-160m, which has 12 layers. When compared to Figures 26 and 27 in Appendix M, for all layers pairs (excluding Layers 0), and for both SVCCA and RSA scores, almost all the p-values of the residual stream layer pairs are between 0% to 1%, indicating that the feature space similarities are statistically meaningful. In particular, Figure 2 shows that Layers 2 and 3, which are middle layers of Pythia-70m, are more similar by SVCCA and RSA scores to middle layers 4 to 7 of Pythia-160m compared to other layers. Figure 20 also shows that Layer 5, the last layer of Pythia-70m, is more similar to later layers of Pythia-160m than to other layers. The number of features after filtering non-concept and many-to-1 features is given in Tables 5 and 6 in Appendix M.
@@
Gemma-2-2B vs Gemma-2-9B. We obtain activations from OpenWebText using 150 samples with a max sequence length of 150, for a total of 22.5k tokens. Figure 4 shows SVCCA and RSA scores after filtering keywords and for 1-1. In particular, the L11 vs L21 comparison achieves an SVCCA score of 0.70 (while mean random pairing has a score of 0.009) and an RSA score of 0.195 (while mean random pairing for has a score of $4.38\times 10^{-4}$). For these experiments, we use 10 runs for mean random pairings at only one layer pair due to the longer run times for larger model pairs.
Llama 3-8B-Instruct vs Llama 3.1-8B . We compare Llama 3-8B-Instruct and Llama 3.1-8B at each model’s Layer 25 residual stream, a later layer in both models. We obtain activations using 100 samples with a max sequence length of 100, for 10000 total tokens. After filtering to keep 7% of Llama 3.1-8B feature pairs, this comparison achieves an SVCCA score of 0.3, comparable to the score of later layers in both Gemma and Pythia, and a mean activation correlation score of 66%.
@@ 4.3 Similarity of Semantically Matched SAE Feature Spaces
We find that for all layers and for all concept categories, Test 2 described in §3 is passed. Thus, we only report specific results for Test 1 in Tables 1 and 2. Overall, in both Pythia and Gemma models and for many concept categories, we find that semantic subspaces are more similar to one another than non-semantic subspaces. More discussion on these results is given in Appendix L.
Rather than just finding that middle layers are more similar to middle layers, we also find that they are the best at representing concepts. This is consistent with work by (Rimsky et al., 2024), which find the best steering vectors in middle layers for steering behaviors such as sycophancy. Other papers find that middle layers tend to provide better directions for eliminating refusal behavior (Arditi et al., 2024) and work well for representing goals and finding steering vectors in policy networks (Mini et al., 2023).
Pythia-70m vs Pythia-160m. We compare every layer of Pythia-70m to every layer of Pythia-160m for several concept categories. While many layer pairs have similar semantic subspaces, middle layers appear to have the semantic subspaces with the highest similarity. Table 1 demonstrates one example of this by comparing the SVCCA score for layer 3 to layer 5 of Pythia-160m, which shows high similarity for semantic subspaces across models, as the concept categories all pass Test 1, having p-values below 0.05. We show the scores for other layer pairs and categories in Figures 30 and 32 in Appendix M, which include RSA scores in Table 9.
@@ 5 Related Work
Feature Universality. Previous works studied universality in terms of components such as features and circuits (Olah and Batson, 2023; Chughtai et al., 2023). Bricken et al. (2023) measure individual SAE feature similarity for two 1-layer toy models; however, this study did not analyze the global properties of feature spaces. Gurnee et al. (2024) find evidence of universal neurons across language models by applying pairwise correlation metrics. Ye et al. (2024) discover "feature families" representing related hierarchical concepts in SAEs across semantically different datasets. Huh et al. (2024) showed that as vision and language models are trained with more parameters and with better methods, their representational spaces converge towards more similar representations. Crosscoders (Lindsey et al., 2024; Mishra-Sharma et al., 2025) and Universal SAEs (Thasarathan et al., 2025) are SAE variations that capture individual features shared across models.
@@
Representational Space Similarity. Previous work has studied neuron activation spaces by utilizing metrics to compare the geometric similarities of representational spaces (Raghu et al., 2017; Wang et al., 2018b; Kornblith et al., 2019; Klabunde et al., 2024, 2023; Kriegeskorte et al., 2008; Sucholutsky et al., 2023), finding that even models with different architectures may share similar representation spaces, hinting at feature universality. However, these techniques have yet to be applied to the feature spaces of SAEs trained on LLMs. Moreover, these techniques compare the similarity of paired input activations. Our work differs as it compares the similarity of paired feature weights. Essentially, previous works do not show that one can match LLMs by features learned by both LLMs. For vision models, Kondapaneni et al. (2025) discover shared and unique visual concepts between different models.
@@
Mechanistic Interpretability. Previous work has made notable progress in neuron interpretation (Foote et al., 2023; Garde et al., 2023) and interpreting the interactions between components (Neo et al., 2024). Other work in mechanistic interpretability focus on circuits analysis (Elhage et al., 2021), as well using steering vectors to control model behaviors (Zou et al., 2023; Turner et al., 2024). These approaches have been combined with SAEs to steer models in the more interpretable SAE feature space (Chalnev et al., 2024). Steering vectors have been found across models, such as vectors that steered a model to refuse harmful instructions across 13 different models (Arditi et al., 2024).
@@
AI Safety. Advanced AI may not be well aligned with the values of their designers (Ngo et al., 2022), due to phenomena such as goal misgeneralization (Shah et al., 2022), deceptive alignment (Greenblatt et al., 2024), and reward tampering (Denison et al., 2024). SAEs can find steerable safety-related features across models, such as refusal features (O’Brien et al., 2024). Finding reliable universal capabilities and control methods, such as in SAEs, may help provide solutions to AI safety issues.
@@ 6 Conclusion
In this study, we take the first steps in investigating the unexplored topic of feature space universality in LLMs via SAEs trained on different models. To achieve these insights, we developed a novel methodology to pair features with high activation correlations, and then assess the global similarity of the resulting subspaces using SVCCA and RSA. Our findings reveal a high degree of SVCCA similarity in SAE feature subspaces across various models, particularly between their middle layers. Furthermore, our research reveals that subspaces of features associated with specific semantic concepts, such as calendar or people tokens, demonstrate remarkably high similarity across different models. This suggests that certain semantic feature subspaces may be universally encoded across varied LLM architectures. However, our results also show that there are parts of SAE feature spaces which do not match, and may suggest that either certain features are not captured universally, or that SAEs have limitations in capturing universal features. We believe that future work can benefit from these discoveries to find better ways on improving methods to capture feature universality. Lastly, while we discover evidence that LLMs learn "weakly universal" features, and that SAEs can capture both these features and these relations, we also find that there is a notable fraction of features idiosyncractic to different SAEs, a finding that is supported by previous work (Leask et al., 2025; Paulo and Belrose, 2025). Thus, our research guides future work to improve upon methods that can better capture universal feature representations, and contributes to ongoing research that aims to improve SAEs (Engels et al., 2024). Our approach may be generalized to compare feature spaces beyond those of SAEs.
Examining feature universality implications for AI safety and control, particularly in identifying and mitigating potentially harmful representations, can contribute to the development of safer AI systems (Barez et al., 2023). Overall, understanding the commonalities and differences in model representations is a crucial step towards a more comprehensive theory of how models encode language.
@@
We perform analysis on only a sample of SAEs for similar LLMs that use the same tokenizer, and focus on SAEs with the same or similar number of features. As we pair features by activation correlation, we do not perform analysis on models that use different tokenizers. Although, for some examples, we obtain good results when pairing featuers only by semantic similarity, these results were not consistent when applied to other examples, perhaps due to the lower pairing quality, as the representational similarity measures depend on the pairing choice. Future work can analyze models whose representations do not depend as much on the tokenizer choice (team et al., 2024).
Our study is also limited by the inherent challenges in interpreting and comparing high-dimensional feature spaces. The methods we used to measure similarity, while robust, may not capture all nuances of feature interactions and representations within the SAEs. Additionally, our analysis is constrained by the computational resources available, which limited our activations dataset sizes, along with the scale and depth of our comparisons. The generalizability of our findings to SAEs trained on more diverse datasets or specialized domain-specific LLMs remains an open question. Lastly, our analysis does not account for potential temporal dynamics in feature representation that may occur at different training stages of LLMs and SAEs.
@@
Michael conceived the study and designed the experiments with help from Fazl. Michael performed the experiments with help from Austin and led the writing of the manuscript. Philip, Ashkan and David provided advice throughout the project. Fazl advised this work and helped significantly with the write-up of the paper.